# Disentangling Representations through Multi-task Learning

### v2

@@ Abstract
Intelligent perception and interaction with the world hinges on internal representations that capture its underlying structure (“disentangled” or “abstract” representations). Disentangled representations serve as world models, isolating latent factors of variation in the world along approximately orthogonal directions, thus facilitating feature-based generalization. We provide experimental and theoretical results guaranteeing the emergence of disentangled representations in agents that optimally solve multi-task evidence aggregation classification tasks, canonical in the cognitive neuroscience literature. The key conceptual finding is that, by producing accurate multi-task classification estimates, a system implicitly represents a set of coordinates specifying a disentangled representation of the underlying latent state of the data it receives. The theory provides conditions for the emergence of these representations in terms of noise, number of tasks, and evidence aggregation time. Surprisingly, the theory also produces closed-form expressions for extracting the disentangled representation from the model’s latent state $\mathbf{Z}(t)$. We experimentally validate these predictions in RNNs trained on multi-task classification, which learn disentangled representations in the form of continuous attractors, leading to zero-shot out-of-distribution (OOD) generalization in predicting latent factors. We demonstrate the robustness of our framework across autoregressive architectures, decision boundary geometries and in tasks requiring classification confidence estimation. We find that transformers are particularly suited for disentangling representations, which might explain their unique world understanding abilities. Overall, our framework puts forth parallel processing as a general principle for the formation of cognitive maps that capture the structure of the world in both biological and artificial systems, and helps explain why ANNs often arrive at human-interpretable concepts, and how they both may acquire exceptional zero-shot generalization capabilities.
@@ 1 Introduction
The ability to construct representations that capture the underlying structure of the world from data, is a hallmark of intelligence. Humans and animals leverage their experiences to construct such faithful representations of the world ("world models"), resulting in a near-effortless ability to generalize to new settings [42, 43]. Modern foundation models also display emergent out-of-distribution (OOD) generalization abilities, in the form of zero- or few-shot learning [10, 61, 35, 58]; however whether artificial systems learn world models remains unclear. Understanding the conditions under which that occurs is bound to lead to better generalizable systems, and explain why artificial systems often converge to human interpretable, aligned representations of the world [71].
@@
A promising direction towards understanding the construction of world models is abstract, or disentangled representations [30, 38, 37]. These two concepts are interrelated yet somewhat distinct (see definitions adapted from Ostojic and Fusi [59] in Section A.1). Shortly, an abstract representation of $x_{1},\dots,x_{n}$ represents each $x_{i}$ linearly and approximately mutually orthogonally. Disentangled representations encode each $x_{i}$ orthogonally, without the necessity of linearity. Both representations preserve the latent structure present in the world in their geometry by isolating factors of variation in the data, which facilitates downstream generalization. When a representation is abstract, a linear decoder (i.e. downstream neuron) trained to discriminate between two categories can readily generalize to stimuli not observed in training, due to the structure of the representation. Furthermore, the more disentangled the representation is, the lower the interference from other variables and hence the better the performance. This corresponds to decomposing a novel stimulus into its familiar features, and performing feature-based generalization. For instance, imagine you are at a grocery store, deciding whether a fruit is ripe or not. If the brain’s internal representation of food attributes (ripeness, caloric content, etc.) is disentangled, then learning to perform this task for bananas would lead to zero-shot generalization to other fruit (e.g. mangos, Figure 1a). Crucially, the visual representation of a mango is high-dimensional, non-linear and noisy, making it particularly challenging to extract a low dimensional latent like "ripeness".
@@
Several brain areas including the amygdala, prefrontal cortex and hippocampus encode variables of interest in an abstract format [67, 5, 8, 56, 16]. This raises the question of under which conditions do such representations emerge in biological and artificial agents alike. Previous work showed that feedforward neural networks develop abstract representations when trained to multitask [37]. However, real-world decisions typically rely on imperfect, noisy information, evolving dynamically over time [9, 40]. To account for this important feature of the world, we train autoregressive models (RNNs, LSTMs, transformers) to multitask canonical neuroscience tasks involving the accumulation of evidence over noisy streams. The tasks tie closely to Bayesian filtering theory, and should be solved by any agent that deals with a noisy world.
@@
The main contributions of the paper are the following:
@@
  • We confirm that RNNs trained to multitask develop abstract representations that zero-shot generalize OOD, when $N_{\text{task}}\geq D$, and orthogonal, disentangled representations for greater $N_{\text{task}}$. The computational substrate of these representations is a 2D continuous attractor [3] storing a ground truth estimate in a product space of the latent factors. In addition, the representations are sparse and mixed, attributes of biological neural networks.
  • We reproduce these findings in GPT-2 models, which generalize better due to them learning disentangled representations already from $N_{\text{task}}\geq D$, confirming their appropriateness for constructing disentangled world models.
  • We demonstrate that our setting is robust to a number of manipulations, including correlated inputs, interleaved learning of tasks and free reaction-time tasks canonical in the cognitive neuroscience literature [9, 40].
  • Finally, we discuss implications for generalizable representation learning in biological and artificial systems, and demonstrate the strong advantage of multi-task learning over previously proposed mechanisms of representation learning in the brain [49].
Despite being framed in the context of canonical cognitive neuroscience tasks, our results are general; they apply to any system aggregating noisy evidence over time.
@@ 1.1 Related work
Disentanglement has long been recognized as a promising strategy for generalization [4] (although note Locatello et al. [47], Montero et al. [55] for a contrarian view), yet most classic work focuses on feedforward architectures [30, 38, 76]. In terms of autoregressive models, Hsu et al. [33], Li and Mandt [46] showed that variational LSTMs disentagle representations of underlying factors in sequential data allowing style transfer; however the underlying representational geometry was not characterised. Other work focuses on fitting RNNs to behavioral data while enforcing disentanglement for interpretability [18, 53]. Work on context-dependent decision making has shown that RNNs re-purpose learned representations in a compositional manner when trained in related tasks [78, 20]; however, the abstractness of the resulting representations was not established. Finally John et al. [36] show that multitasking results in disentanglement, however unlike us they directly enforce latent factor separation through their adversarial optimization objectives. Our approach is most closely related to weakly supervised disentanglement, without comparing across samples [68].
Our work relates to previous work on linear identifiability. Roeder et al. [64] show that representations of models trained on the same distribution must be linear transformations of each other; yet we go beyond their results to show that abstract representations are guaranteed to emerge under moderate conditions, irrespectively of the dimensionality of the input and model architecture. Lachapelle et al. [41] proved that disentangled representations emerge in feedforward architectures from multitask learning in sparse tasks when a sparsity regularization constraint is placed on the predictors; we place no such constraints and still uncover disentangled representations.
Previous neuroscience-inspired work showed that multitasking feedforward networks learn abstract representations, as quantified by regression generalization [37]. We expand upon these findings in several ways. First, we extend the framework to autoregressive architectures (RNNs, LSTMs, transformers) that can update their representations as further information arrives. Second, we prove theorems that guarantee the emergence of abstract representations in any optimal multitask classifier if the number of tasks exceeds the input dimensionality $D$, and showcase disentanglement in our trained networks. Third, we rigorously analyze the role of noise in forming disentangled representations, extending the noise-free regime studied in Johnston and Fusi [37]. Finally, we explore a range of values for $D$, providing experimental validation of our theory.
@@ 2 Problem formulation
We study the evidence aggregation multi-task classification paradigm shown in Figure 1b. An agent with latent state $\mathbf{Z}(t)$ receives noisy, non-linearly mapped observations $\{f\big{(}\mathbf{X}(t)\big{)}\}_{t=1}^{T}$ where each $\mathbf{X}(t)=\mathbf{x}^{*}+\sigma\,\mathcal{N}(0,I_{D})$ is a noisy measurement of unknown ground truth vector $\mathbf{x}^{*}\in\mathbb{R}^{D}(x_{i}^{*}\sim\textit{Uniform}(-0.5,0.5))$, with $\mathcal{N}$ being Gaussian noise. The noisy measurements are transformed by an injective observation map function $f$, which can be non-linear and high dimensional, representing the wide range of sensory transformations found in real-world scenarios. The agent is tasked with simultaneously solving $N_{\textit{task}}$ classification problems by aggregating information over time, each defined by a random linear decision boundary [^1] Due to the observation map $f$, the tasks may appear non-linear from the perspective of the multi-task classification agent. in the ground truth space $\mathbb{R}^{D}$ i.e.
@@ 3 Theoretical results
$\hat{\mathbf{Y}}(t)$ $=g(\mathbf{Z}(t))$ (3)
The notion of optimality allows us to make precise statements about the informational content of the agent’s internal state $\mathbf{Z}(t)$. Let $\mathbf{C}\in\mathbb{R}^{N_{\textit{task}}\times D}$ be a matrix where each row is a decision boundary normal vector. Then
@@
    any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ must encode a finite-sample, maximum likelihood estimate $\mu(t)$ of the ground truth evidence variable $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$, and
@@
    if the activation function is sigmoid-like, $\mu(t)$ will be linearly decodable from $\mathbf{Z}(t)$, thus implying that $\mathbf{Z}(t)$ contains an abstract representation of $\mu(t)$ [59].
@@
    Furthermore, the representation is guaranteed to be disentangled (orthogonal) as $N_{\textit{task}}\gg D$ for random decision boundaries.
Specifically, $\mu(t)$ is the maximum likelihood estimate (MLE) of $\mathbf{x}^{*}$ given observations $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$. A closed-form expression for extracting $\mu(t)$ from $\mathbf{Z}(t)$ is given by:
@@
The key conceptual insight in the proof of Theorem 3.1 is that each of the multi-task classification probability estimates $\hat{Y}_{i}(t)$ represents an estimated projection distance between the MLE $\mu(t)$ and the given classification boundary $(\mathbf{c}_{i},b_{i})$. Once distances to classification boundaries are recovered, $\mu(t)$ can be inferred if the $N_{\text{task}}$ classification boundaries span the $D$-dimensional space of $\mathbf{x}^{*}$. Note that classification boundary distances are only guaranteed to be recoverable when there is non-zero noise $\sigma>0$ such that $\hat{Y}_{i}(t)$ does not saturate to 1 or 0, and thus still carries useful decision boundary distance information (see Lemma B.3). In fact, Equations 4 and 5 do not hold for $\sigma\to 0$, as they were derived by means of Bayesian estimation which assumes the presence of noise.
The linear approximation for decoding $\mu(t)$ from $\mathbf{Z}(t)$ in Equation 5 is enabled by the remarkable similarity between sigmoid functions and the Gaussian CDF $\Phi$. The sigmoid-like structure of $\Phi$ suggests many similar activations $g$ (e.g., softmax) would exhibit approximate linear decodability. For more exotic element-wise activation functions $g$, Equation 4 suggests $\mu(t)$ will be represented in $g(\mathbf{Z}(t))$ in the column space of $(\mathbf{C^{\top}C})^{-1}\mathbf{C}^{\top}$ potentially with some element-wise non-linearity.
Robustness to other noise distributions
Although we focus on Gaussian noise for our theoretical derivations, we note that a wide range of alternative noise distributions are permissible, including anisotropic Gaussian noise (Lemma B.10). The key property enabling us to recover distances from classification probabilities is the invertibility of the marginal posterior CDFs along the decision boundary normals. This property is not exclusive to Gaussian noise but is shared by a broader class of noise distributions, including but not limited to elliptical distributions (e.g., multivariate t-distributions, Laplace distributions), exponential power distributions, and other symmetric and unimodal distributions with invertible marginals. Therefore, the proof of Theorem 3.1 generalizes to any noise distribution satisfying the conditions outlined above (see Appendix B.4 for conditions and derivations). In fact, our experiments show robustness to anisotropic and autocorrelated noise, and mismatch between noise CDF and activation function (Section A.8).
@@ 4 Methods
where $\mathbf{W}_{\textit{rec}}$ is the recurrent weight matrix, $\mathbf{W}_{\textit{in}}$ is the matrix carrying the input vector $\mathbf{x}_{\textit{in}}$, $\mathbf{b}$ is a unit-specific bias vector, $\tau$ is the neuronal time constant, $\left[.\right]_{+}$ is the ReLU applied element-wise and time dependencies have been dropped for brevity. We discretize Equation 6 using the forward Euler method for $T=20$ timesteps of duration $\Delta t=\tau=100\,\textrm{ms}$, which we find to be stable. The RNN’s output $\hat{\mathbf{y}}(t)\in\mathbb{R}^{N_{\textit{task}}}$ is given by $\hat{\mathbf{y}}(t)=g(\mathbf{W}_{\textit{out}}\,\mathbf{z}(t))$, where $\mathbf{W}_{\textit{out}}$ is a readout matrix and $g=\text{sigmoid}$ the output activation function applied elementwise.
Figure 2: Data generation and architecture. (a) The task is to simultaneously report whether the ground truth $\mathbf{x}^{*}$ lies above ($1$) or below (00) a number of classification lines. (b) Models (RNN depicted) are trained to report the outcome of all the binary classifications in a at the end of the trial (indicated by the fixation input turning 0).
The encoder $f$ is a 3-layer MLP with hidden dimensions $100,100,40$ and ReLU non-linearities, and it is randomly initialized and kept fixed during training as it represents a static mapping from latents to observations (observation map). An additional fixation input is directly passed to the hidden layer. It is 1 during the trial and turns 0 at the end of the trial, indicating that the network should report its decisions (Figure 2b). The fixation input precludes the RNN from learning a specific timing in its response. We refer to this kind of tasks as fixed reaction-time (RT). The network is trained with a cross-entropy loss and Adam default settings, except learning rate $\eta_{0}=10^{-3}$, to produce the target outputs $\mathbf{y}(\mathbf{x}^{*})$. By minimizing loss across trials, the network is incentivized to estimate $\hat{\mathbf{Y}}(t)=\Pr\{y_{i}(\mathbf{x}^{*})=1\}$. Table S1 summarizes all hyperparameters and their values, which are shared across all architectures.

### v3

@@ Abstract
Intelligent perception and interaction with the world hinges on internal representations that capture its underlying structure (“disentangled” or “abstract” representations). Disentangled representations serve as world models, isolating latent factors of variation in the world along approximately orthogonal directions, thus facilitating feature-based generalization. We provide experimental and theoretical results guaranteeing the emergence of disentangled representations in agents that optimally solve multi-task evidence accumulation classification tasks, canonical in the neuroscience literature. The key conceptual finding is that, by producing accurate multi-task classification estimates, a system implicitly represents a set of coordinates specifying a disentangled representation of the underlying latent state of the data it receives. The theory provides conditions for the emergence of these representations in terms of noise, number of tasks, and evidence accumulation time, when the classification boundaries are affine in the latent space. Surprisingly, the theory also produces closed-form expressions for extracting the disentangled representation from the model’s latent state $\mathbf{Z}(t)$. We experimentally validate these predictions in RNNs trained on multi-task classification, which learn disentangled representations in the form of continuous attractors, leading to zero-shot out-of-distribution (OOD) generalization in predicting latent factors. We demonstrate the robustness of our framework across autoregressive architectures, decision boundary geometries and in tasks requiring classification confidence estimation. We find that transformers are particularly suited for disentangling representations, which might explain their unique world understanding abilities. Overall, our framework establishes a formal link between competence at multiple tasks and the formation of disentangled, interpretable world models in both biological and artificial systems, and helps explain why ANNs often arrive at human-interpretable concepts, and how they both may acquire exceptional zero-shot generalization capabilities.
@@ 1 Introduction
The ability to construct representations that capture the underlying structure of the world from data, is a hallmark of intelligence. Humans and animals leverage their experiences to construct such faithful representations of the world (”world models”, ”cognitive maps”), resulting in a near-effortless ability to generalize to new settings (Lake et al., 2015; 2016). Modern foundation models also display emergent out-of-distribution (OOD) generalization abilities, in the form of zero- or few-shot learning (Brown et al., 2020; Pham et al., 2021; Jia et al., 2021; Oquab et al., 2023); however whether artificial systems learn world models remains unclear. Understanding the conditions under which that occurs is bound to lead to better generalizable systems, and explain why artificial systems often converge to human interpretable, aligned representations of the world (Templeton et al., 2024).
@@
A promising direction towards understanding the construction of world models is abstract, or disentangled representations (Higgins et al., 2017; Kim & Mnih, 2018; Johnston & Fusi, 2023). These two concepts are interrelated yet somewhat distinct (see definitions adapted from Ostojic & Fusi (2024) in Section A.1). Shortly, an abstract representation of $x_{1},\dots,x_{n}$ represents each $x_{i}$ linearly and approximately mutually orthogonally. Disentangled representations encode each $x_{i}$ orthogonally, without the necessity of linearity. Both representations preserve the latent structure present in the world in their geometry by isolating factors of variation in the data, which facilitates downstream generalization. When a representation is abstract, a linear decoder (i.e. downstream neuron) trained to discriminate between two categories can readily generalize to stimuli not observed in training, due to the structure of the representation. Furthermore, the more disentangled the representation is, the lower the interference from other variables and hence the better the performance. This corresponds to decomposing a novel stimulus into its familiar features, and performing feature-based generalization. For instance, imagine you are at a grocery store, deciding whether a fruit is ripe or not. If the brain’s internal representation of food attributes (ripeness, caloric content, etc.) is disentangled, then learning to perform this task for bananas would lead to zero-shot generalization to other fruit (e.g. mangos, Figure 1a). Crucially, the visual representation of a mango is high-dimensional, non-linear and noisy, making it particularly challenging to extract a low dimensional latent like ”ripeness”.
@@
Several brain areas including the amygdala, prefrontal cortex and hippocampus encode variables of interest in an abstract format (Saez et al., 2015; Bernardi et al., 2020; Boyle et al., 2022; Nogueira et al., 2023; Courellis et al., 2024). This raises the question of under which conditions do such representations emerge in biological and artificial agents alike. Previous work showed that feedforward neural networks develop abstract representations when trained to multitask (Johnston & Fusi, 2023). However, real-world decisions typically rely on imperfect, noisy information, evolving dynamically over time (Britten et al., 1992; Krajbich et al., 2010). To account for this important feature of the world, we train autoregressive models (RNNs, LSTMs, transformers) to multitask canonical neuroscience tasks involving the accumulation of evidence over noisy streams. The tasks tie closely to Bayesian filtering theory, and should be solved by any agent that deals with a noisy world.
@@
Our main contributions are the following:
@@
  • We confirm that RNNs trained to multitask develop abstract representations that zero-shot generalize OOD, when $N_{\text{task}}\geq D$, and orthogonal, disentangled representations for greater $N_{\text{task}}$. The computational substrate of these representations is a 2D continuous attractor (Amari, 1977) storing a ground truth estimate in a product space of the latent factors. In addition, the representations are sparse and mixed, attributes of biological neural networks.
  • We reproduce these findings in GPT-2 transformers, which generalize better due to them learning disentangled representations already from $N_{\text{task}}\geq D$, confirming their appropriateness for constructing disentangled world models.
  • We demonstrate that our setting is robust to a number of manipulations, including correlated inputs, interleaved learning of tasks and free reaction-time tasks canonical in the cognitive neuroscience literature (Britten et al., 1992; Krajbich et al., 2010).
  • Finally, we discuss implications for generalizable representation learning in biological and artificial systems, and demonstrate the strong advantage of multi-task learning over previously proposed mechanisms of representation learning in the brain (Mante et al., 2013).
Although framed in the context of canonical neuroscience tasks, our results are general; they apply to any system aggregating noisy evidence. While our experiments focus on supervised multi-task learning for tractability, the theory only assumes competence at multiple tasks, thus enabling alternative methods of acquiring such competence, such as self-supervised or unsupervised pre-training.
@@ 1.1 Related work
Disentanglement has long been recognized as a promising strategy for generalization (Bengio et al., 2012) (although note Locatello et al. (2019); Montero et al. (2020) for a contrarian view), yet most classic work focuses on feedforward architectures (Higgins et al., 2017; Kim & Mnih, 2018; Whittington et al., 2022; Maziarka et al., 2023). In autoregressive models, Hsu et al. (2017); Li & Mandt (2018) showed that variational LSTMs disentagle representations of underlying factors in sequential data allowing style transfer; however the underlying representational geometry was not characterised. Other work focuses on fitting RNNs to behavioral data while enforcing disentanglement for interpretability (Dezfouli et al., 2019; Miller et al., 2023). Work on context-dependent decision making has shown that RNNs re-purpose learned representations in a compositional manner when trained in related tasks (Yang et al., 2019; Driscoll et al., 2022); however, the abstractness of the resulting representations was not established. Finally John et al. (2018) show that multitasking results in disentanglement, however unlike us they directly enforce latent factor separation through their adversarial optimization objectives. Our approach is most closely related to weakly supervised disentanglement, without comparing across samples (Shu et al., 2019).
Our work relates to previous work on linear identifiability. Roeder et al. (2021) show that representations of models trained on the same distribution must be linear transformations of each other; yet we go beyond their results to show that abstract representations are guaranteed to emerge under moderate conditions, irrespectively of the dimensionality of the input and model architecture. Lachapelle et al. (2023) proved that disentangled representations emerge in feedforward architectures from multitask learning in sparse tasks when a sparsity regularization constraint is placed on the predictors; we place no such constraints and still uncover disentangled representations.
Previous neuroscience-inspired work showed that multitasking feedforward networks learn abstract representations, as quantified by regression generalization (Johnston & Fusi, 2023). We expand upon these findings in several ways. First, we extend the framework to autoregressive architectures (RNNs, LSTMs, transformers) that can update their representations as further information arrives. Second, we prove theorems that guarantee the emergence of abstract representations in any optimal multitask classifier if the number of tasks exceeds the input dimensionality $D$, and showcase disentanglement in our trained networks. Third, we rigorously analyze the role of noise in forming disentangled representations, extending the noise-free regime studied in Johnston & Fusi (2023). Finally, we explore a range of values for $D$, providing experimental validation of our theory.
@@ 2 Problem formulation
We study the evidence accumulation multi-task classification paradigm shown in Figure 1b. An agent with latent state $\mathbf{Z}(t)$ receives noisy, non-linearly mapped observations $\{f\big{(}\mathbf{X}(t)\big{)}\}_{t=1}^{T}$ where each $\mathbf{X}(t)=\mathbf{x}^{*}+\sigma\,\mathcal{N}(0,I_{D})$ is a noisy measurement of unknown ground truth vector $\mathbf{x}^{*}\in\mathbb{R}^{D}(x_{i}^{*}\sim\textit{Uniform}(-0.5,0.5))$, $\mathcal{N}$ being Gaussian noise. The noisy measurements are transformed by an injective observation map $f$, which can be non-linear and high dimensional, representing the wide range of sensory transformations found in real-world scenarios. The agent is tasked with simultaneously solving $N_{\textit{task}}$ classification problems by accumulating information over time (a canonical neuroscience task (Britten et al., 1992)), each defined by a random linear decision boundary [^1] Due to the observation map $f$, the tasks may appear non-linear from the perspective of the multi-task classification agent. in the ground truth space $\mathbb{R}^{D}$ i.e.
@@ 3 Theoretical results
The notion of optimality allows us to make precise statements about the informational content of the agent’s internal state since $\hat{\mathbf{Y}}(t)=g(\mathbf{Z}(t))$. Let $\mathbf{C}\in\mathbb{R}^{N_{\textit{task}}\times D}$ be a matrix where each row is a decision boundary normal vector. Then
@@
    Any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ must encode a finite-sample, maximum likelihood estimate $\mu(t)$ of the ground truth evidence variable $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$.
@@
    If the activation function is sigmoid-like, $\mu(t)$ will be linearly decodable from $\mathbf{Z}(t)$, thus implying that $\mathbf{Z}(t)$ contains an abstract representation of $\mu(t)$ (Ostojic & Fusi, 2024).
@@
    The representation is guaranteed to be disentangled (orthogonal) as $N_{\textit{task}}\gg D$ for random decision boundaries.
Specifically, $\mu(t)$ is the maximum likelihood estimate (MLE) of $\mathbf{x}^{*}$ given observations $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$. A closed-form expression for extracting $\mu(t)$ from $\mathbf{Z}(t)$ if $N_{\textit{task}}\geq D$ is:
@@
The key conceptual insight in the proof of Theorem 3.1 is that each of the multi-task classification probability estimates $\hat{Y}_{i}(t)$ represents an estimated projection distance between the MLE $\mu(t)$ and the given classification boundary $(\mathbf{c}_{i},b_{i})$. Once distances to classification boundaries are recovered, $\mu(t)$ can be inferred if the $N_{\text{task}}$ classification boundaries span the $D$-dimensional space of $\mathbf{x}^{*}$.
Robustness of results
While Theorem 3.1 applies to optimal multi-task classifiers, Corollary B.7 shows that a sub-optimal multi-classifier with zero-mean independent errors will represent $\tilde{\mu}(t)$ in state $\mathbf{Z}(t)$ (Equation 3) with residual errors w.r.t. optimal $\mu(t)$ expected to decrease at a rate of approximately $\mathcal{O}(1/\sqrt{N_{task}})$. See Appendix B.4, B.9 for extensions of the theory to anisotropic and non-Gaussian noise distributions (Elliptical, t-distribution, Laplace distributions). The linear approximation for decoding $\mu(t)$ from $\mathbf{Z}(t)$ in Equation 4 is enabled by the remarkable similarity between sigmoid functions and the Gaussian CDF $\Phi$ (Corollary B.8). The sigmoid-like structure of $\Phi$ suggests many similar activations $g$ (e.g., softmax) would exhibit approximate linear decodability.
More general decision boundaries
Decision boundaries $y_{i}$ on latents $\mathbf{x}^{*}$ may appear non-linear in the image of observation map $f$, but Theorem 3.1 applies to linear boundaries $y_{i}$ on latent space (Eq.  1). Our results extend naturally to smooth manifold decision boundaries through local linearization when the manifold $y_{i}$’s reach $\tau_{i}$222“Reach”: maximum distance at which each point on a manifold has a unique closest point on the manifold is much larger than the noise scale $\sigma$. Intriguingly, classification boundary distances are only guaranteed to be recoverable when there is non-zero noise $\sigma>0$ such that $\hat{Y}_{i}(t)$ does not saturate to 1 or 0, and thus still carries useful decision boundary distance information (see Lemma B.3)  [^3] In fact, Equations 3 and 4 do not hold for $\sigma\to 0$, as they were derived by means of Bayesian estimation which assumes the presence of noise. . An intriguing open question is what conditions on manifolds $\{y_{i}\}_{i=1}^{N_{\text{task}}}$ are necessary and sufficient to preserve the decodability of $\mu(t)$. We leave a complete characterization of representation learning with multiple manifold decision boundaries for future work.
@@ 4 Methods
Figure 2: Data generation and architecture. (a) For each trial, we sample a ground truth vector $\mathbf{x}^{*}$, and add IID noise to arrive at $\mathbf{X}(t)$. The task is to report whether $\mathbf{x}^{*}$ lies above ($1$) or below (00) each of the classification lines (color matches corresponding boolean variable in $y$), given the noisy and non-linearly transformed samples $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$. (b) Models (RNN depicted) are trained to report the outcome of all the binary classifications in a at the end of the trial (indicated by the fixation input turning 0).
@@
where $\mathbf{W}_{\textit{rec}}$ is the recurrent weight matrix, $\mathbf{W}_{\textit{in}}$ is the matrix carrying the input vector $\mathbf{x}_{\textit{in}}$, $\mathbf{b}$ is a unit-specific bias vector, $\tau$ is the neuronal time constant, $\left[.\right]_{+}$ is the ReLU applied element-wise and time dependencies have been dropped for brevity. We discretize Equation 5 using the forward Euler method for $T=20$ timesteps of duration $\Delta t=\tau=100\,\textrm{ms}$, which we find to be stable. The RNN’s output $\hat{\mathbf{y}}(t)\in\mathbb{R}^{N_{\textit{task}}}$ is given by $\hat{\mathbf{y}}(t)=g(\mathbf{W}_{\textit{out}}\,\mathbf{z}(t))$, where $\mathbf{W}_{\textit{out}}$ is a readout matrix and $g=\text{sigmoid}$ the output activation function applied elementwise. The encoder $f$ is a 3-layer MLP with hidden dimensions $100,100,40$ and ReLU non-linearities, and it is randomly initialized and kept fixed during training as it represents a static mapping from latents to observations (observation map). An additional fixation input is directly passed to the hidden layer. It is 1 during the trial and turns 0 at the end of the trial, indicating that the network should report its decisions (Figure 2b). The fixation input is concatenated with $f(\mathbf{X}(t))$ to form $\mathbf{x}_{\textit{in}}$, and it precludes the RNN from learning a specific timing in its response. We refer to this kind of tasks as fixed reaction-time (RT). The network is trained with a cross-entropy loss and Adam default settings, except learning rate $\eta_{0}=10^{-3}$, to produce the target outputs $\mathbf{y}(\mathbf{x}^{*})$. By minimizing loss across trials, the network is incentivized to estimate $\hat{\mathbf{Y}}(t)=\Pr\{y_{i}(\mathbf{x}^{*})=1\}$. Table S1 summarizes all hyperparameters and their values, which are shared across all architectures.